### v2

# Pre-Processing Matters: Similarity-based Simulated Annealing Approach for Fast Parameter Tuning in WSI Classification

Abstract
Pre-processing whole slide images (WSIs) can impact classification performance. Our study shows that using fixed hyper-parameters for pre-processing out-of-domain WSIs can significantly degrade performance. Therefore, it is critical to search domain-specific hyper-parameters during inference. However, searching for an optimal parameter set is time-consuming. To overcome this, we propose SSAPT, a novel Similarity-based Simulated Annealing approach for fast parameter tuning to enhance inference performance on out-of-domain data. The proposed SSAPT achieves 5% to 50% improvement in accuracy with ×$\times 5$ times faster parameter searching speed on average.
Introduction
Following the success of early CPATH applications, dataset sizes have increased, prompting more multicentric efforts to address variability in staining, image quality, scanning characteristics, and tissue preparation across different labs. This has highlighted a known issue where CPATH algorithms perform best on data they were trained on but less well on data from other sources. Generalization continues to pose a major challenge in this field, as significant differences in clinical variables between tissue source sites can adversely affect the performance of histopathological tasks (Van der Laak, Litjens, and Ciompi 2021; Dehkharghanian et al. 2023; Howard et al. 2021).
State-of-the-art MIL models have shown promising improvements in WSI classification on out-of-domain data by training on large, diverse datasets (Zhang et al. 2022; Campanella et al. 2019; Lu et al. 2021; Shao et al. 2021; Li, Li, and Eliceiri 2021; Tang et al. 2023). Out-of-domain classification occurs when a model is trained on one dataset and tested on another from a different domain. For instance, a model might be trained using Camelyon16 data (Litjens et al. 2018) and then evaluated using data from various centers in Camelyon17. However, these studies typically only report average accuracy, neglecting the variability in performance across different centers. As indicated in Table 1, the performance of a model trained on Camelyon16 can vary significantly at specific centers within Camelyon17.
Figure 1: Different pre-processing methods for Whole Slide Images (WSIs) generate distinct features, which in turn affect inference performance. Retaining excessive background and discarding too much tissue can result in poor inference outcomes.
We discovered that unsatisfactory performance is often due to using the same inference hyper-parameter across different centers. In real-world scenarios, we apply the fixed pre-processing parameters to extract features and evaluate performance on new sub-datasets. However, this approach typically results in poor out-of-domain performance (Howard et al. 2021). For example, MIL models (Lu et al. 2021; Cui et al. 2023; Li, Li, and Eliceiri 2021) have extremely low accuracy at specific centers within TCGA and Camelyon17 datasets. Conversely, tailoring optimal pre-processing parameters for each center significantly improves out-of-domain performance (as illustrated in Figure 2).
Determining optimal pre-processing parameters for histopathological tasks is challenging due to the vast size of Whole Slide Images (WSIs) and the significant computational resources needed. Traditional grid search methods for datasets like Camelyon or TCGA are inefficient, taking several hours to evaluate one parameter group. The whole hyper-parameter space contains hundreds of thousands of possible combinations. Therefore, quicker hyper-parameter search techniques are essential for histopathological tasks.
Figure 2: The performance of out-of-domain inference varies with preprocessing parameters across various MIL models and datasets. Consequently, we suggest that each specific center within the dataset should adopt its own preprocessing parameters to maximize performance. The original method involves all centers using fixed default preprocessing hyperparameters, whereas the optimal method allows each center to use its own specific preprocessing hyperparameters determined by our proposed SSAPT.
We have observed that varying pre-processing methods can significantly impact model performance during both inference and training. This paper focuses on how pre-processing parameters influence inference performance, particularly with out-of-domain data. This paper proposes SSAPT, Similarity-based Simulated Annealing for Hyper-parameter Tuning, a method that improved the accuracy of inferring out-of-domain data by 5% to 50% across various MIL models and datasets. Previous studies, such as those cited in (Choi et al. 2024), only explored hyper-parameter optimization in terms of per-layer learning rates and loss weight coefficients , without tailoring their approaches to histopathology applications. The key contributions of this paper are:
  • We have observed that varying pre-processing parameters significantly impact feature extraction and, consequently, model performance—particularly in out-of-domain inference;
  • We present SSAPT, Similarity-based Simulated Annealing for fast and effective parameter tuning. This algorithm enhances inference performance at Camelyon 17’s center 1, boosting accuracy from 0.512 to 0.847;
  • We expand the proposed SSAPT to include other MIL models using various public datasets, achieving improvements in accuracy ranging from 5% to 50% for out-of-domain data across multiple MIL models and datasets;
  • The proposed SSAPT is the first fast hyper-parameter search specifically designed for histopathological tasks.
Related work
MIL models. DSMIL and CLAM employ instance-loss for training, with DSMIL featuring an instance-level classifier. Bayes-MIL focuses on patch-level uncertainty. TransMIL (Shao et al. 2021) treats patches as tokens to utilize transformer architecture. MHIM (Tang et al. 2023) employs a masked hard instance mining strategy. All these models are designed at the instance/patch level, which leads to poor out-of-domain performance. In contrast, ABMIL (Ilse, Tomczak, and Welling 2018) operates solely at the bag level, resulting in lower accuracy in in-domain performance across various datasets but achieving greater robustness. However, it is impossible to completely avoid instance-level design in developing new MIL models for histopathology.
Pre-processing of WSI in histopathology. Variations in tissue processing techniques, including chemical fixation or freezing, dehydration, embedding, and staining, can alter the visual characteristics of the tissue slide in ways that are both non-uniform and non-linear. These changes occur across different tissue types and laboratories (Asif et al. 2023; Yagi 2011), consequently impact the performance in deep learning. Salvi et al.; Gurcan et al. summize the pre-processing impact for WSI analysis on deep learning frameworks. Tellez et al. quantifies the effects of data augmentation and stain color normalization. (Lee et al. 2022)
There are several research works that specifically explore the issue of dataset bias in CPath. Pocevičiūtė et al. quantifies the impact of domain shift in attention-based MIL and point out that MIL performance is worse than what is reported using in-domain test data. Dehkharghanian et al.; Howard et al. discusses bias on histopathological data. Important clinical variables have been shown to be significantly different between tissue source sites, affecting the generalization issue for histopathological tasks.
Hyper-parameters search and optimization in machine learning. Falkner, Klein, and Hutter propose a Bayes-optimization with Hyperband strategy and evaluate it in small datasets. Choi et al. search hyper-parameters in per-layer learning rates and loss weight coefficients. These previous works are not designed for histopathology.
Motivation and Observation
Table 1: Performance of Out-of-Domain Inference in Specific Challenging Case. The performance is reported as the average of Accuracy, AUC, Precision, Recall, and F-score metrics, computed over the 10-fold models.
Model In-Domain (C16) Out-of-Domain (C17 Center 1)
Acc AUC Accuracy AUC Precision Recall F-score
Min Max Min Max Min Max Min Max Min Max
ABMIL 0.901 0.941 0.84 0.92 0.833 0.896 0.788 0.966 0.743 0.829 0.765 0.875
DSMIL 0.925 0.954 0.64 0.89 0.721 0.93 0.491 1.000 0.429 0.943 0.556 0.831
CLAM 0.901 0.946 0.512 0.847 0.708 0.833 0.486 0.839 0.743 1.000 0.648 0.788
TransMIL 0.892 0.935 0.73 0.84 0.861 0.914 0.730 0.823 0.753 0.844 0.724 0.830
Bayes-MIL 0.883 0.916 0.35 0.80 0.819 0.881 0.350 0.759 0.629 1.000 0.519 0.737
MHIM-DSMIL 0.925 0.965 0.77 0.87 0.850 0.921 0.607 1.000 0.600 0.971 0.710 0.800
Public datasets often combine data from various domains (Torralba and Efros 2011; Pocevičiūtė et al. 2023), obscuring true performance metrics within specific areas. Studies (Dehkharghanian et al. 2023; Howard et al. 2021; Hosseini et al. 2024) show that domain-specific variations in digital histology can affect the accuracy and bias of deep learning models. In practice, these variations necessitate frequent retraining of models due to changing data domains, which is impractical due to high computational costs. Thus, improving model performance without additional training is essential.
Pre-processing Impacts Inference Performance for Different Domains.
We found there exists significant variability in inference performance on out-of-domain data across various methods (refer to Table 1). For instance, the CLAM model trained on the Camelyon16 dataset achieves an overall inference accuracy of 0.86 on the Camelyon17 dataset but drops to 0.512 at Center 1 of Camelyon17, indicating inconsistent performance. We have also tested our approach in other centers and datasets, as detailed in the Appendix.
We employ a grid search to fully understand how pre-processing parameters influence outcomes. We discovered that modifying these parameters affects out-of-domain (OOD) inference accuracy differently across centers. The default pre-processing setting yields varying results depending on the center as shown in Fig 3, where identical hyper-parameters may perform well in one center but poorly in another. Similar variations are observed across other datasets and MIL models (additional details provided in Appendix). This inconsistency stems from instance/patch-level design choices. We conduct experiments to examine how adjusting bag loss and instance loss weights for DSMIL and CLAM impacts OOD inference performance. Comprehensive details about these instance-level design experiments are available in the Appendix.
Inefficient Manual Pre-Processing Parameter Search.
It’s beneficial to optimize parameter performance, but finding the best settings is time-consuming; it takes approximately 7 hours to extract features and evaluate a set of parameters on an RTX3080 (details in Appendix). Therefore, we need more efficient parameter search methods. Common strategies like grid search require excessive computational resources and time, rendering them impractical for this scenario.
Based on these observations, we want to explore three questions.
  • Can modifying only the preprocessing improve inference performance on out-of-domain data?
  • How can we search for parameters more cost-effectively to enhance inference performance?
  • Do these observations apply to other state-of-the-art (SOTA) models?
Therefore, we introduce our approach in fast hyper-paramater search in pre-processing parameters of histopathology.
Figure 3: The inference performance of out-of-domain data varies with preprocessing parameters across various MIL models and datasets.
Fast Hyper-Parameter Search for Inference Performance in Out-Of-Domain Data
In this section, we formalize hyperparameter optimization and present our SSAPT algorithm step by step. The proposed SSAPT algorithm includes PSNR-based mechanism for limiting the current searching space in each iteration and Simulated Annealing to avoid local optima.
Definition of Problem
A good choice of hyperparameters can significantly improve out-of-domain inference performance, yet optimal hyperparameters vary, depending on the datasets and the models. Formally, we denote the fast hyper-parameter search algorithm as $H_{opt}=SSAPT(H,D_{val}^{n})$ and its hyper-parameters as $H\in\Phi$, where $\Phi$ is the hyper-parameter space. We denote the training and validation datasets as $D_{train}$ and $D_{val}$, respectively. The model $M$ is trained in $D_{train}$, where all WSIs are using fixed hyper-parameter $H_{train}$. We denote the next inferring processing as $y=Infer(H_{i},D_{val}^{n})$. Hyperparameter optimization finds hyperparameters that maximize some performance metric $y$. We formalize hyperparameter optimization as
$H_{n}^{*}=\underset{H\in\Phi}{\operatorname{arg\,max}}\,\text{Infer}(SSAPT(H,D_{val}^{n}),D_{val}^{n})$ (1)
General Assumptions
We assume access to two types of datasets: (1) a large dataset $D_{train}$ for training to produce the foundation model subsequently used to test out-of-domain performance. (2) Multiple subdataset $\{D_{val}=D_{val}^{1},D_{val}^{2},\cdots,D_{val}^{n}\}$. The validation sub-dataset $D_{val}$ is much smaller than $D_{train}$ and is used only for optimization at the outer level, not for fine-tuning. Note that these test distributions are distinct and are never shown to the model before the final evaluation.
Let $M$ denote a pre-trained foundation model trained in a large dataset $D_{train}$ with the pre-processing parameter $H_{train}$. Note that the model $M$ is trained by k-fold, StratifiedKFold, or Monte Carlo cross-validation, so $D_{train}$ actually contains a sub-dataset of training, validation, and testing while training.
Now, there are many out-of-domain datasets $\{D_{val}=D_{val}^{1},D_{val}^{2},\cdots,D_{val}^{n}\}$. They are pre-processed by hyper-parameter $H_{1},H_{2},\cdots,H_{n}$ to obtain the features $F_{val}^{1},F_{val}^{2},\cdots,F_{val}^{n}$ for inferring. The out-of-domain performance of model $M$ would be tested by the features $F_{val}^{n}$ that are preprocessed by the hyper-parameter $H_{n}$. We adapt our algorithms as $H_{n}^{*}=SSAPT(H_{n},D_{val}^{n}$) to search the optimal hyper-parameters $H_{n}$ for out-of-domain dataset $D_{val}^{n}$.
In our experiments, the performance metric $y$ is accuracy, due to the unbalanced datasets in histopathological tasks. For instance, some centers contain only one category of WSI in TCGA-NSCLC, resulting in the situation that AUC and F-score cannot be calculated. Furthermore, extremely low accuracy is reported in some challenging cases, leading to the situation that accuracy is lower than 0.5, but the AUC is still higher than 0.8 (as shown in Table 1).
Pre-Processing Problem Modeling
We define the task of identifying the optimal hyper-parameter for enhanced out-of-domain performance as an optimization problem, as described in Equation 1. In practice, this involves repeatedly applying different features within the same foundational model using various hyperparameters $H_{n}^{i}$. This process helps us determine the most effective final hyper-parameters $H_{n}$ for each validation center $D_{val}^{n}$. We assess the effectiveness of each $H_{n}^{i}$ by evaluating its performance on the OOD validation set $D_{val}^{n}$.
Next we consider the optimization problem for histopathological tasks. The optimization of $H_{n}$ is actually the optimization of a combination of pre-processing parameters, not a single variable. Previous pre-processing of WSIs encompasses a series of standard procedures. Initially, tissue segmentation is executed automatically on each slide at a reduced magnification. This involves generating a binary mask for the tissue regions by applying a binary threshold to the saturation channel of the downsampled image, following its conversion from RGB to HSV color space. Morphological closing and median blurring are employed for smoothing the contours of the detected tissue and in minimizing artifacts. Subsequently, the approximate contours of both the tissue and the tissue cavities are assessed and filtered based on their area, culminating in the creation of the final segmentation mask. We complete set of pre-processing hyperparameters for optimization in histopathological tasks is thus $H_{n}=[x_{1},x_{2},x_{3},x_{4},x_{5},x_{6}]$, where each $x_{i}$ has its own hyper-parameter space.
Optimization of Fast Hyper-Parameter Search
Good Pre-Processed WSIs show High Structural Similarity.
For single-domain inference performance, our experiments demonstrate that pre-processing parameters yielding accurate predictions produce images with similar structures (as illustrated in Figure 5). Therefore, we only evaluate the pre-processed Whole Slide Images (WSIs) that closely resemble the optimally pre-processed WSI. To quantitatively assess differences between two sets of pre-processing results, we use Peak Signal-to-Noise Ratio (PSNR) to measure the similarity of images generated during patch creation. As depicted in Figure 1, our workflow includes creating patches, extracting features, training models, and testing inference data. During patch generation, thumbnails are created for visualizing the preprocessing effects. Thus, we can determine the similarity between these two sets of pre-processed results by calculating PSNR for these thumbnails.
PSNR
Peak Signal-to-Noise Ratio (PSNR) is a common measurement in image processing and digital signal processing. A higher PSNR value indicates greater similarity between two images. We calculate the PSNR by comparing thumbnails generated during preprocessing. For example, the total PSNR of thumbnails produced using the second-best preprocessing parameters should be higher than that of any other set, as illustrated in Figure 5.
[…]
In our motivation experiments, which explore the impact of pre-processing parameters on out-of-domain performance, we have already processed sufficient Whole Slide Images (WSIs) with inference results. We established the similarity threshold $\tau$ based on previous experimental analyses. For our proposed SSAPT method, we use ${H_{train}}$ as the default parameter. Using the current optimal parameter $H^{*}$, we generate a new parameter ${H_{new}}$, create patches, and calculate their PSNR against the optimal thumbnails $I^{*}$. If the PSNR exceeds threshold $\tau$, we proceed to more resource-intensive tasks such as feature extraction in histopathological analysis and subsequent inference using an existing model. Each new parameter set ${H_{new}}$ is added to our historical dataset $R=\{(H_{1},y_{1}),...,(H_{i},y_{i})\}$. If ${H_{new}}$ demonstrates improved performance, there is a probability $p$ (as per Eq.(3)) that ${H_{*}}$ will be updated with ${H_{new}}$. This process will be further discussed in the following subsection.
Similarity-based Simulated Annealing for Parameter Tuning
Algorithm 1 Similarity based Simulated Annealing for Parameter Tuning (SSAPT)
Input: Default parameter ${H_{train}}$ with result set $\mathcal{R}$, Hyper-parameters space $\Phi$, initial temperature $T_{0}$, minimum temperature $T_{\text{min}}$, cooling rate $\alpha$, similarity threshold $\tau$, function $Infer$ to evaluate performance, function $S$ to calculate similarity, $p$ is the acceptance probability function
Output: Optimal parameters $H^{*}$
1:  Initialize $H^{*}$ with fixed default parameter ${H_{train}}$
2:  Generate pre-processed thumbnail $I^{*}$ from $H^{*}$
3:  Generate feature $F^{*}$ from WSIs with $H^{*}$
4:  $y^{*}\leftarrow Infer(F^{*})$
5:  $\mathcal{R}\leftarrow(H^{*},y^{*})$
6:  $T\leftarrow T_{0}$
7:  while $T>T_{\text{min}}$ do
8:     Select $H_{\text{new}}$ by perturbing $H^{*}$ slightly from $\Phi$
9:     if  $H_{\text{new}}$ not in $\mathcal{R}$ then
10:        Generate thumbnail $I_{\text{new}}$ with $H_{\text{new}}$
11:        $\Delta s\leftarrow S(I_{\text{new}},I^{*})-\tau$
12:        if $\Delta s>0$  then
13:           Extract feature $F_{\text{new}}$      $\triangleright$ Expensive step
14:           $y_{\text{new}}\leftarrow Infer(F_{\text{new}})$
15:           $\mathcal{R}\leftarrow\mathcal{R}\cup(H_{\text{new}},y_{\text{new}})$
16:           if $y_{\text{new}}>y^{*}$  then
17:              $H^{*},I^{*},y^{*}\leftarrow{H_{\text{new}}},I_{\text{new}},y_{\text{new}}\text{ with $p$}$
18:           end if
19:        end if
20:        $T\leftarrow\alpha T$
21:     end if
22:  end while
23:  return  $H^{*}$
To avoid falling into local optima, we integrate the Simulated Annealing algorithm (Kirkpatrick, Gelatt Jr, and Vecchi 1983) into our method to identify optimal parameters. We detail SSAPT in Algorithm 1. Our experimental results indicate that creating patches with thumbnails and bypassing most parameter feature extraction processes allows us to refine pre-processing parameters for enhanced inferential performance.
Our algorithm is designed to optimize pre-processing parameters effectively, ensuring high performance while accelerating the overall process. The Simulated Annealing algorithm prevents convergence at local optima by exploring multiple random states near the current solution, akin to simulated annealing itself. Occasionally, we enhance solutions by leveraging the current temperature to explore nearby states post-annealing. This approach introduces randomness through annealing rules, significantly increasing the likelihood of achieving an optimal solution tailored for histopathological tasks.
During each iteration, we aim to match the pre-processed results closely with the current optimum. We use a similarity threshold $\tau$, related to PSNR as a filter mechanism. After patch creation, features are extracted and analyzed only from pre-processed Whole Slide Images (WSIs) that exhibit a PSNR higher than $\tau$. WSIs with lower PSNR compared to the optimal pre-processed WSI are excluded from further processing. If superior inference performance emerges, updates are made for $H^{*},I^{*},y^{*}$ based on acceptance probability $p$:
$p(\Delta s)=\exp(\Delta s/T)$ (3)
Experiments
Dataset Center ABMIL DSMIL CLAM Bayes-MIL MHIM-DSMIL
Default Ours Default Ours Default Ours Default Ours Default Ours
Camelyon17 C0 0.93 0.97 0.77 0.96 0.921 0.95 0.83 0.94 0.87 0.96
C1 0.87 0.92 0.85 0.89 0.512 0.846 0.35 0.8 0.81 0.87
C4 0.88 0.93 0.92 0.96 0.874 0.92 0.85 0.92 0.91 0.94
TCGA-COAD CM 0.75 0.778 0.75 0.778 0.75 0.889 0.861 0.889 0.75 0.801
D5 0.709 0.774 0.806 0.871 0.742 0.871 0.709 0.806 0.839 0.871
DM 0.783 0.826 0.869 0.869 0.739 0.739 0.783 0.826 0.826 0.87
G4 0.963 0.963 0.963 0.963 0.889 0.926 0.926 1 0.925 0.963
TCGA-BRCA A7 0.902 0.961 0.921 0.941 0.941 0.961 0.882 0.882 0.923 0.961
AC 0.775 0.8 0.725 0.8 0.75 0.8 0.775 0.775 0.75 0.775
AR 0.781 0.8 0.765 0.781 0.813 0.828 0.813 0.828 0.75 0.766
C8 0.907 0.93 0.581 0.813 0.907 0.93 0.86 0.884 0.907 0.93
Table 2: Experiments in out-of-domain inference performance. We compare the accuracy obtained by default hyper-parameters with the optimal hyper-parameter searched by our SSAPT algorithm.
Experiments Setup
Our experiment utilizes the pre-processing pipeline from CLAM. The datasets employed in this study include Camelyon 16, Camelyon 17 (Litjens et al. 2018), TCGA-NSCLC, TCGA-BRCA, and TCGA-COAD (Weinstein et al. 2013). For Camelyon16 and Camelyon17, we focus on normal-tumor classification; for all TCGA datasets, we address subtype tasks.
In our experiments, we standardize seeds and model hyperparameters such as learning rate and loss weight across similar tasks within the same model framework. Additional experimental details are provided in the Appendix. We employ K-fold, StratifiedKFold or k-fold Monte Carlo cross-validation methods (with k set to 10) to train models and assess out-of-domain performance. All results presented in Table 2 and Table 3 represent average accuracies calculated over these 10-fold models.
Classification Accuracy Validation of SSAPT
For the Camelyon datasets, we trained our model using 270 whole slide images (WSIs) from Camelyon16 and tested inference performance across various centers in Camelyon17 to simulate out-of-domain scenarios. It’s important to note that Camelyon16 data was collected from UMCU and RUMC, while Camelyon17 includes data from CWZ, RST, UMCU, RUMC, and LPON. Consequently, only centers 0 (CWZ), 1 (RST), and 4 (LPON) of Camelyon17 represent out-of-domain situations.
Figure 4: The running time of pre-processing parameters tuning for 100 times. Our algorithm can save time by skipping a large number of feature extraction.
We compared the inference performance using features extracted by default parameters (not the worst) and those optimized through our SSAPT algorithm. As shown in Table 2, the optimal hyper-parameters identified by our SSAPT algorithm generally surpass the default settings across various models and datasets. In specific cases, such as with CLAM and BayesMIL at Center 1, performance improved dramatically from below 0.5 to up to 0.8 simply by adjusting the hyper-parameters. Even for robust models like ABMIL, modifying parameters proved effective.
For TCGA datasets, we divide the entire dataset according to TCGA Tissue Source Site Codes. We select several centers with a sufficient number of WSIs to form the testing sub-dataset, while the remaining data constitutes the training dataset. The out-of-domain inference performance also improves (as shown in Table 2), indicating that our issue is general and our SSAPT provides a universal solution. Additionally, we conduct experiments on a larger TCGA-NSCLC dataset that includes more extensive out-of-domain testing (as detailed in Table 3).
Dataset Center ABMIL CLAM Bayes-MIL
Def. Ours Def. Ours Def. Ours
TCGA- NSCLC 21 0.722 0.778 0.556 0.778 0.556 0.667
22 0.865 0.892 0.784 0.838 0.838 0.838
43 0.750 0.800 0.600 0.750 0.700 0.800
49 0.595 0.643 0.690 0.714 0.714 0.762
55 0.495 0.516 0.659 0.758 0.275 0.330
77 0.932 0.932 0.909 0.932 0.750 0.800
97 1.000 1.000 0.913 1.000 0.869 0.957
Table 3: Comparison of inference accuracy in TCGA-NSCLC. We evaluate the default hyper-parameter with optimal hyper-parameter searched by our SSAPT Algorithm in a larger dataset TCGA-NSCLC with more centers.
Efficiency Validation
Figure 4 shows the running time for tuning preprocessing parameters 100 times. Our Simulated Annealing algorithm for Parameter Tuning uses PSNR to compare each pre-processed WSI with the best one so far, selecting only those with high similarity for further feature extraction. WSIs with low PSNR are skipped, streamlining the process and ensuring that only promising candidates are evaluated further. This approach prevents our SSAPT from getting stuck in local optima and saves time by reducing unnecessary feature extractions.
Ablation Study
PSNR
Without the PSNR-based strategy, the algorithm performs a random search. It randomly generates pre-processing parameters, creates patches, extracts features, and assesses performance in each iteration without skipping feature extraction (as shown in Table 4).
Figure 5: The evidence of PSNR-based mechanism works in our fast hyper-parameter search in histopathological tasks.
We also assess FDD-Simulated Annealing. It’s important to note that the Fréchet domain distance (FDD) in this part of the ablation study serves a different role compared to its use in the Discussion section. Here, FDD replaces PSNR for comparison purposes. Both FDD and PSNR-based mechanisms in this ablation study are employed to constrain the search space for optimal hyper-parameters, rather than searching across the entire hyper-parameter space.
Table 4: Importance of the PSNR. Results are based on optimizing 100 times in the center1 of Camelyon17. SA stands for Simulated Annealing, Lat. stands for Latency and Mem. stands for Memory.
Strategy Acc Lat. (min) Mem. (GB)
None (Baseline) 0.847 9600 1250
FDD-SA 0.846 5350 650
PSNR-SA (Ours) 0.846 1770 170
Simulated Annealing
Simulated Annealing prevents optimization from getting stuck in local optima. As illustrated in Table 5, search algorithms lacking Simulated Annealing repeatedly test hyper-parameters that yield a higher PSNR than the current optimal solution.
Table 5: Importance of Simulated Annealing (SA). Results based on optimizing 100 times in center1 of Camelyon17.
Strategy Accuracy Latency (min)
Without SA 0.829 4119
With SA 0.846 1770
Discussion
Quantifying the Initial Features
We demonstrate that significant differences exist in the initial features extracted by various pre-processing parameters, which are used as inputs for training models or making inferences. These differences arise not only from distribution shifts but also from variations in the number of patches extracted from each WSI, leading to partial information loss. For the same WSI, we compare features generated under different hyper-parameter settings: $H_{opt}$, $H_{good}$, and $H_{random}$. As shown in Figures 1 and  6, features are similar when their inference performance is comparable. However, Figure 6 (Right) highlights distinct differences between $H_{opt}$ and $H_{random}$ within the t-SNE space, underscoring how patch count variations can exacerbate feature representation discrepancies and affect out-of-distribution (OOD) inference performance.
Figure 6: The visualization of compasion between features. Left: The comparison of the feature obtaining high accuracy with the optimal feature. Right: The comparison of the feature extracted by random hyper-parameters with the optimal one. Note that these features are extracted from the same WSIs.
Evaluating the Influence of Pre-processing on Model Representations
We utilize Fréchet domain distance (FDD) (Pocevičiūtė et al. 2023) to compare features from the penultimate layer and assess how pre-processing parameters influence the model’s representations. For each specific out-of-domain testing dataset $D_{val}^{i}$, WSIs undergo pre-processing using $H_{i}$ to generate feature $F_{i}$. This feature is then input into the model for inference. We extract the penultimate layer feature $P_{i}$ before it reaches the final classifier. The same process applies when using optimal pre-processing features identified by our SSAPT algorithm, denoted as $H_{opt}$, resulting in $F_{opt}$. For WSIs processed with $H_{i}$, a higher accuracy correlates with a lower FDD compared to those processed with optimal preprocessing ($H_{opt}$), as illustrated in Figure 7.
Figure 7: Quantifying the Features Impact by Model Visibility.
SSAPT is designed for histopathological tasks
Previous hyper-parameter optimization stategies (Falkner, Klein, and Hutter 2018; Choi et al. 2024) have been developed with a focus on general machine learning models and evaluated on datasets characterized by a large number of images but relatively small image sizes. However, the unique characteristics of WSIs should not be ignored. WSIs are giga-pixel images that demand significantly higher computational resources. Despite advancements in hardware, memory constraints still pose a significant challenge in the digital processing of WSIs in current research (Cifci et al. 2023; Echle et al. 2021). Additionally, nature images have very different data distribution compared to the whole slide image (Raghu et al. 2019). Therefore, designing hyper-parameter optimization stategy for CPath is necessary. The proposed SSAPT algorithm is the first hyper-parameter optimization stategy designed for histopatholocial tasks.
Conclusion
Different pre-processing parameters significantly impact feature extraction and model performance in histopathological tasks. In this paper, we propose the Similarity-based Simulated Annealing (SSAPT) algorithm for fast parameter tuning, which enhances inference performance on out-of-domain data. SSAPT achieves a 5% to 50% accuracy improvement on the Camelyon and multiple TCGA datasets, offering faster hyper-parameter search by reducing feature extraction steps based on WSI characteristics.

### v3

@@
-Pre-Processing Matters: Similarity-based Simulated Annealing Approach for Fast Parameter Tuning in WSI Classification
+BAHOP: Similarity-based Basin Hopping for A fast hyper-parameter search in WSI classification
 Abstract
@@
 However, searching for an optimal parameter set is time-consuming.
-To overcome this, we propose SSAPT, a novel Similarity-based Simulated Annealing approach for fast parameter tuning to enhance inference performance on out-of-domain data.
-The proposed SSAPT achieves 5% to 50% improvement in accuracy with ×$\times 5$ times faster parameter searching speed on average.
-Introduction
-Following the success of early CPATH applications, dataset sizes have increased, prompting more multicentric efforts to address variability in staining, image quality, scanning characteristics, and tissue preparation across different labs.
-This has highlighted a known issue where CPATH algorithms perform best on data they were trained on but less well on data from other sources.
-Generalization continues to pose a major challenge in this field, as significant differences in clinical variables between tissue source sites can adversely affect the performance of histopathological tasks (Van der Laak, Litjens, and Ciompi 2021; Dehkharghanian et al.
-2023; Howard et al.
-2021).
-State-of-the-art MIL models have shown promising improvements in WSI classification on out-of-domain data by training on large, diverse datasets (Zhang et al.
-2022; Campanella et al.
-2019; Lu et al.
-2021; Shao et al.
-2021; Li, Li, and Eliceiri 2021; Tang et al.
-2023).
+To overcome this, we propose BAHOP, a novel Similarity-based Basin Hopping optimization for fast parameter tuning to enhance inference performance on out-of-domain data.
+The proposed BAHOP achieves 5% to 30% improvement in accuracy with ×$\times 5$ times faster on average.
+1 Introduction
+Following the success of early computational pathology applications, dataset sizes have increased, prompting more multicentric efforts to address variability in staining, image quality, scanning characteristics, and tissue preparation in different laboratories.
+This has highlighted a known issue where computational pathology algorithms perform best on data on which they were trained but less well on data from other sources.
+Generalization continues to pose a major challenge in this field, as significant differences in clinical variables between tissue source sites can adversely affect the performance of histopathological tasks [26, 6, 11].
+State-of-the-art multiple instance learning (MIL) models have shown promising improvements in whole slide image (WSI) classification on out-of-domain (OOD) data by training on large, diverse datasets [31, 2, 17, 21, 15, 23].
 Out-of-domain classification occurs when a model is trained on one dataset and tested on another from a different domain.
-For instance, a model might be trained using Camelyon16 data (Litjens et al.
-2018) and then evaluated using data from various centers in Camelyon17.
-However, these studies typically only report average accuracy, neglecting the variability in performance across different centers.
-As indicated in Table 1, the performance of a model trained on Camelyon16 can vary significantly at specific centers within Camelyon17.
-Figure 1: Different pre-processing methods for Whole Slide Images (WSIs) generate distinct features, which in turn affect inference performance.
-Retaining excessive background and discarding too much tissue can result in poor inference outcomes.
-We discovered that unsatisfactory performance is often due to using the same inference hyper-parameter across different centers.
-In real-world scenarios, we apply the fixed pre-processing parameters to extract features and evaluate performance on new sub-datasets.
-However, this approach typically results in poor out-of-domain performance (Howard et al.
-2021).
-For example, MIL models (Lu et al.
-2021; Cui et al.
-2023; Li, Li, and Eliceiri 2021) have extremely low accuracy at specific centers within TCGA and Camelyon17 datasets.
+For instance, a model might be trained using Camelyon16 data [16] and then evaluated using data from various Camelyon17 centers.
+As indicated in Table 1, the performance of a model trained in Camelyon16 can vary significantly at specific centers within Camelyon17.
+Figure 1: The performance of out-of-domain inference varies with preprocessing parameters across various MIL models and datasets.
+Consequently, we suggest that each specific center within the dataset should adopt its own preprocessing parameters to maximize performance.
+The original method involves all centers using fixed default preprocessing hyperparameters, whereas the optimal method allows each center to use its own specific preprocessing hyperparameters determined by our proposed BAHOP.
+Figure 2: PSNR-based Basin Hopping hyper-parameter optimization for out-of-domain inference in WSIs.
+The hyper-parameter optimization for the pre-processing task is a non-convex optimization that contains many large local optima.
+Our BAHOP is developed for fast search, and the jumping range of each iteration is controlled by the PSNR threshold.
+We discovered that one factor that leads to unsatisfactory performance is the same inference hyper-parameter across different centers.
+We evaluate the fixed pre-processing parameters’ performance on several sub-datasets.
+The experiment result shows that fixed pre-processing parameters typically result in poor out-of-domain performance [11].
+For example, MIL models [17, 5, 15] have extremely low accuracy at specific centers within the TCGA [29] and Camelyon17 [16] datasets using fixed parameters.
 Conversely, tailoring optimal pre-processing parameters for each center significantly improves out-of-domain performance (as illustrated in Figure 1).
-Determining optimal pre-processing parameters for histopathological tasks is challenging due to the vast size of Whole Slide Images (WSIs) and the significant computational resources needed.
-Traditional grid search methods for datasets like Camelyon or TCGA are inefficient, taking several hours to evaluate one parameter group.
+Determining optimal pre-processing parameters for histopathological tasks is challenging.
+The parameter search space is over $10^{5}$ due to the huge size of giga-pixel Whole Slide Images (WSIs) and the significant computational resources needed.
+Previous studies [3] only explore hyper-parameter optimization in learning rates and loss weight coefficients without tailoring their approaches to histopathology applications.
+When adopting hyper-parameter tuning methods, traditional grid search methods for datasets like Camelyon or TCGA are inefficient, taking several hours to evaluate one parameter group.
 The whole hyper-parameter space contains hundreds of thousands of possible combinations.
 Therefore, quicker hyper-parameter search techniques are essential for histopathological tasks.
-Figure 2: The performance of out-of-domain inference varies with preprocessing parameters across various MIL models and datasets.
-Consequently, we suggest that each specific center within the dataset should adopt its own preprocessing parameters to maximize performance.
-The original method involves all centers using fixed default preprocessing hyperparameters, whereas the optimal method allows each center to use its own specific preprocessing hyperparameters determined by our proposed SSAPT.
-We have observed that varying pre-processing methods can significantly impact model performance during both inference and training.
-This paper focuses on how pre-processing parameters influence inference performance, particularly with out-of-domain data.
-This paper proposes SSAPT, Similarity-based Simulated Annealing for Hyper-parameter Tuning, a method that improved the accuracy of inferring out-of-domain data by 5% to 50% across various MIL models and datasets.
-Previous studies, such as those cited in (Choi et al.
-2024), only explored hyper-parameter optimization in terms of per-layer learning rates and loss weight coefficients , without tailoring their approaches to histopathology applications.
+This paper focuses on how pre-processing parameters influence inference performance, particularly with out-of-domain data, and proposes an efficient parameter search method specifically designed for WSI classification.
+We propose a Similarity-based Basin Hopping (BAHOP) for Hyper-parameter tuning and improving the accuracy of inferring out-of-domain data across various MIL models and datasets (as shown in Fig.
+1).
 The key contributions of this paper are:
   • We have observed that varying pre-processing parameters significantly impact feature extraction and, consequently, model performance—particularly in out-of-domain inference;
-  • We present SSAPT, Similarity-based Simulated Annealing for fast and effective parameter tuning.
-This algorithm enhances inference performance at Camelyon 17’s center 1, boosting accuracy from 0.512 to 0.847;
-  • We expand the proposed SSAPT to include other MIL models using various public datasets, achieving improvements in accuracy ranging from 5% to 50% for out-of-domain data across multiple MIL models and datasets;
-  • The proposed SSAPT is the first fast hyper-parameter search specifically designed for histopathological tasks.
-Related work
-MIL models.
-DSMIL and CLAM employ instance-loss for training, with DSMIL featuring an instance-level classifier.
-Bayes-MIL focuses on patch-level uncertainty.
-TransMIL (Shao et al.
-2021) treats patches as tokens to utilize transformer architecture.
-MHIM (Tang et al.
-2023) employs a masked hard instance mining strategy.
-All these models are designed at the instance/patch level, which leads to poor out-of-domain performance.
-In contrast, ABMIL (Ilse, Tomczak, and Welling 2018) operates solely at the bag level, resulting in lower accuracy in in-domain performance across various datasets but achieving greater robustness.
-However, it is impossible to completely avoid instance-level design in developing new MIL models for histopathology.
-Pre-processing of WSI in histopathology.
-Variations in tissue processing techniques, including chemical fixation or freezing, dehydration, embedding, and staining, can alter the visual characteristics of the tissue slide in ways that are both non-uniform and non-linear.
-These changes occur across different tissue types and laboratories (Asif et al.
-2023; Yagi 2011), consequently impact the performance in deep learning.
-Salvi et al.; Gurcan et al.
-summize the pre-processing impact for WSI analysis on deep learning frameworks.
-Tellez et al.
-quantifies the effects of data augmentation and stain color normalization.
-(Lee et al.
-2022)
-There are several research works that specifically explore the issue of dataset bias in CPath.
-Pocevičiūtė et al.
-quantifies the impact of domain shift in attention-based MIL and point out that MIL performance is worse than what is reported using in-domain test data.
-Dehkharghanian et al.; Howard et al.
-discusses bias on histopathological data.
-Important clinical variables have been shown to be significantly different between tissue source sites, affecting the generalization issue for histopathological tasks.
-Hyper-parameters search and optimization in machine learning.
-Falkner, Klein, and Hutter propose a Bayes-optimization with Hyperband strategy and evaluate it in small datasets.
-Choi et al.
-search hyper-parameters in per-layer learning rates and loss weight coefficients.
-These previous works are not designed for histopathology.
-Motivation and Observation
+  • We present BAHOP, Similarity-based Basin Hopping for fast and effective parameter tuning.
+This algorithm enhances inference performance at Camelyon 17’s center 1, boosting accuracy from 0.512 to 0.846;
+  • We expand the proposed BAHOP to include other MIL models using various public datasets, achieving improvements in accuracy ranging from 5% to 30% for out-of-domain data across multiple MIL models and datasets;
+  • The proposed BAHOP is the first fast hyper-parameter search designed explicitly for histopathological tasks.
 Table 1: Performance of Out-of-Domain Inference in Specific Challenging Case.
@@
 MHIM-DSMIL 0.925 0.965 0.77 0.87 0.850 0.921 0.607 1.000 0.600 0.971 0.710 0.800
-Public datasets often combine data from various domains (Torralba and Efros 2011; Pocevičiūtė et al.
-2023), obscuring true performance metrics within specific areas.
-Studies (Dehkharghanian et al.
-2023; Howard et al.
-2021; Hosseini et al.
-2024) show that domain-specific variations in digital histology can affect the accuracy and bias of deep learning models.
+2 Related Work
+Pre-processing of WSI in histopathology.
+Variations in tissue processing techniques, including chemical fixation or freezing, dehydration, embedding, and staining, can alter the visual characteristics of the tissue slide in ways that are both non-uniform and non-linear.
+These changes occur across different tissue types and laboratories [1, 30], consequently, impacting the performance of deep learning.
+Salvi et al.
+[20], Gurcan et al.
+[9], Wang et al.
+[28] summarize the pre-processing impact for WSI analysis on deep learning frameworks.
+Tellez et al.
+quantifies the effects of data augmentation and stain color normalization [14].
+There are several research works that specifically explore the issue of dataset bias in computational pathology.
+Pocevičiūtė et al.
+[18] quantifies the impact of domain shift in attention-based MIL and points out that MIL performance is worse than what is reported using in-domain test data.
+Several works [6, 11] discuss bias in histopathological data.
+Important clinical variables have been shown to be significantly different between tissue source sites, affecting the generalization issue for histopathological tasks.
+Hyper-parameters search and optimization in machine learning.
+[8] propose a Bayes-optimization with Hyperband strategy and evaluate it in small datasets.
+[3] search hyper-parameters in per-layer learning rates and loss weight coefficients.
+These previous works are not designed for histopathology.
+Hyper-parameter optimization in histopathological tasks is different than optimization problems in a linear function.
+The optimization in linear function or small machine learning model with a small dataset can be optimized in many iterations in a short time.
+However, the optimization in the machine learning model for histopathological tasks is totally different.
+WSIs are giga-pixel images that demand significantly higher computational resources.
+Despite advancements in hardware, memory constraints still pose a significant challenge in the digital processing of WSIs in current research [4, 7].
+Additionally, nature images have very different data distribution compared to the whole slide image [19].
+Therefore, designing a hyper-parameter optimization strategy for computational pathology is necessary.
+The proposed BAHOP algorithm is the first hyper-parameter optimization strategy designed for histopathological tasks.
+3 Motivation and Observation
+Public datasets often combine data from various domains [25, 18], obscuring true performance metrics within specific areas.
+Studies [6, 11, 10] show that domain-specific variations in digital histology can affect the accuracy and bias of deep learning models.
 In practice, these variations necessitate frequent retraining of models due to changing data domains, which is impractical due to high computational costs.
 Thus, improving model performance without additional training is essential.
-Pre-processing Impacts Inference Performance for Different Domains.
+3.1 Pre-processing Impacts Inference Performance for Different Domains.
+Figure 3: The inference performance of out-of-domain data varies with preprocessing parameters across various MIL models and datasets.
+Figure 4: All the heatmaps circled by red boxes correspond to the default hyper-parameter and predict wrong, while heatmaps circled by blue boxes correspond to optimal hyperparameters with correct predictions.
+Fig.A: Hyper-parameter optimization starts from the default parameter, which is the same as the pre-trained model.
+The default pre-processed WSI drops many patches that get a high attention score in optimal pre-processed WSI.
+Fig.B, the dataset is TCGA-NSCLC where dropping tissue can improve accuracy.
+Fig.C: Some hyper-parameters drop the region of interest(ROI) during pre-processing.
 We found there exists significant variability in inference performance on out-of-domain data across various methods (refer to Table 1).
-For instance, the CLAM model trained on the Camelyon16 dataset achieves an overall inference accuracy of 0.86 on the Camelyon17 dataset but drops to 0.512 at Center 1 of Camelyon17, indicating inconsistent performance.
-We have also tested our approach in other centers and datasets, as detailed in the Appendix.
-We employ a grid search to fully understand how pre-processing parameters influence outcomes.
-We discovered that modifying these parameters affects out-of-domain (OOD) inference accuracy differently across centers.
-The default pre-processing setting yields varying results depending on the center as shown in Fig 3, where identical hyper-parameters may perform well in one center but poorly in another.
-Similar variations are observed across other datasets and MIL models (additional details provided in Appendix).
-This inconsistency stems from instance/patch-level design choices.
-We conduct experiments to examine how adjusting bag loss and instance loss weights for DSMIL and CLAM impacts OOD inference performance.
-Comprehensive details about these instance-level design experiments are available in the Appendix.
-Inefficient Manual Pre-Processing Parameter Search.
-It’s beneficial to optimize parameter performance, but finding the best settings is time-consuming; it takes approximately 7 hours to extract features and evaluate a set of parameters on an RTX3080 (details in Appendix).
-Therefore, we need more efficient parameter search methods.
-Common strategies like grid search require excessive computational resources and time, rendering them impractical for this scenario.
+We employ a grid search to fully understand how pre-processing parameters influence outcomes in center 1 of Camelyon 17 in this motivation experiment.
+For instance, the CLAM [17] model trained on the Camelyon16 dataset achieves an overall inference accuracy of 0.86 on the entire Camelyon17 dataset but drops to 0.512 at Center 1 of Camelyon17, indicating inconsistent performance.
+We also evaluate our approach in other centers across different datasets, as detailed in Fig 3 and Appendix.
+Similar variations are observed across other datasets and MIL models, and the out-of-domain inference is extremely poor in some cases.
+What’s more, the default pre-processing setting yields varying results depending on the center as shown in Fig 3, where identical hyper-parameters may perform well in one center but poorly in another.
+Observation 1
+The effection of the pre-processing parameters varies across different datasets.
+As shown in Fig.
+4 and the Appendix, the detection of foreground contours, controlled by several hyperparameters, often results in the exclusion of large areas, which are treated as holes.
+For instance, adipocyte cells are frequently excluded.
+However, some tumor cells surrounded by fibers, adipocyte cells, or lipid droplets may also be filtered out due to preprocessing parameters and discarded.
+In Fig.
+4(a) and (c), excessive patch removal leads to incorrect predictions, while in Fig.
+4(b), it improves prediction accuracy when applied to a different dataset.
+This variability makes it difficult to quantify the impact of specific patches, especially since the number of patches for a single WSI in a MIL model can exceed $10^{5}$.
+While some studies, such as Javed et al.
+[13], address this issue, there is no consensus on how specific patches influence machine learning models in computational pathology.
+An adaptive algorithm is needed to address this challenge in histopathological tasks.
+Observation 2
+Pre-processing optimization is inherently a non-convex problem with numerous large local optima (as shown in Fig.
+2).
+Our objective is to achieve an optimal result with reduced time and computational costs, where this optimum does not need to be a global one.
+Rather than identifying specific cell types within tissue areas, we approach this histopathological task from a machine learning perspective.
+The pre-processing of WSIs focuses on segmenting tissue regions and discarding patches identified as holes or backgrounds.
+These patches are selected and filtered based on pre-defined parameters, while most tissue areas containing cells are retained.
+Consequently, the variations in inference performance caused by multiple interacting hyperparameters can be framed as a hyperparameter optimization problem.
+This problem can be effectively addressed using an adaptive algorithm tailored for this histopathological task.
+3.2 Inefficient Manual Pre-Processing Parameter Search.
+In real-world scenarios, retraining machine learning models for every new dataset is often impractical.
+While optimizing parameter performance is advantageous, finding the optimal settings is highly time-consuming.
+For instance, evaluating a single set of parameters, including feature extraction, takes approximately 7 hours on an RTX3080 (details in Appendix).
+Traditional optimization techniques, such as tuning learning rates, batch sizes, or loss weight coefficients, cannot be directly applied to histopathological tasks.
+Consequently, more efficient parameter search methods are required.
+Common strategies like grid search demand excessive computational resources and time, rendering them unsuitable for histopathological applications.
 Based on these observations, we want to explore three questions.
-  • Can modifying only the preprocessing improve inference performance on out-of-domain data?
+  • Can we improve inference performance on out-of-domain data just by modifying the preprocessing parameters?
   • How can we search for parameters more cost-effectively to enhance inference performance?
-  • Do these observations apply to other state-of-the-art (SOTA) models?
-Therefore, we introduce our approach in fast hyper-paramater search in pre-processing parameters of histopathology.
-Figure 3: The inference performance of out-of-domain data varies with preprocessing parameters across various MIL models and datasets.
-Fast Hyper-Parameter Search for Inference Performance in Out-Of-Domain Data
-In this section, we formalize hyperparameter optimization and present our SSAPT algorithm step by step.
-The proposed SSAPT algorithm includes PSNR-based mechanism for limiting the current searching space in each iteration and Simulated Annealing to avoid local optima.
-Definition of Problem
+  • Do these observations apply to other multiple instance learning (MIL) models?
+Therefore, we introduce our approach in a fast hyper-parameter search in pre-processing parameters of histopathology.
+4 Fast Hyper-Parameter Search for Inference Performance in Out-Of-Domain Data
+In this section, we formalize hyperparameter optimization and present our BAHOP algorithm step by step.
+The proposed BAHOP algorithm includes PSNR-based mechanism for limiting the current searching space in each iteration and Basin Hopping to avoid local optima.
+4.1 Definition of Problem
 A good choice of hyperparameters can significantly improve out-of-domain inference performance, yet optimal hyperparameters vary, depending on the datasets and the models.
-Formally, we denote the fast hyper-parameter search algorithm as $H_{opt}=SSAPT(H,D_{val}^{n})$ and its hyper-parameters as $H\in\Phi$, where $\Phi$ is the hyper-parameter space.
-We denote the training and validation datasets as $D_{train}$ and $D_{val}$, respectively.
-The model $M$ is trained in $D_{train}$, where all WSIs are using fixed hyper-parameter $H_{train}$.
-We denote the next inferring processing as $y=Infer(H_{i},D_{val}^{n})$.
-Hyperparameter optimization finds hyperparameters that maximize some performance metric $y$.
-We formalize hyperparameter optimization as
-$H_{n}^{*}=\underset{H\in\Phi}{\operatorname{arg\,max}}\,\text{Infer}(SSAPT(H,D_{val}^{n}),D_{val}^{n})$ (1)
-General Assumptions
 We assume access to two types of datasets: (1) a large dataset $D_{train}$ for training to produce the foundation model subsequently used to test out-of-domain performance.
@@
 Let $M$ denote a pre-trained foundation model trained in a large dataset $D_{train}$ with the pre-processing parameter $H_{train}$.
-Note that the model $M$ is trained by k-fold, StratifiedKFold, or Monte Carlo cross-validation, so $D_{train}$ actually contains a sub-dataset of training, validation, and testing while training.
 Now, there are many out-of-domain datasets $\{D_{val}=D_{val}^{1},D_{val}^{2},\cdots,D_{val}^{n}\}$.
 They are pre-processed by hyper-parameter $H_{1},H_{2},\cdots,H_{n}$ to obtain the features $F_{val}^{1},F_{val}^{2},\cdots,F_{val}^{n}$ for inferring.
-The out-of-domain performance of model $M$ would be tested by the features $F_{val}^{n}$ that are preprocessed by the hyper-parameter $H_{n}$.
-We adapt our algorithms as $H_{n}^{*}=SSAPT(H_{n},D_{val}^{n}$) to search the optimal hyper-parameters $H_{n}$ for out-of-domain dataset $D_{val}^{n}$.
-In our experiments, the performance metric $y$ is accuracy, due to the unbalanced datasets in histopathological tasks.
-For instance, some centers contain only one category of WSI in TCGA-NSCLC, resulting in the situation that AUC and F-score cannot be calculated.
-Furthermore, extremely low accuracy is reported in some challenging cases, leading to the situation that accuracy is lower than 0.5, but the AUC is still higher than 0.8 (as shown in Table 1).
-Pre-Processing Problem Modeling
-We define the task of identifying the optimal hyper-parameter for enhanced out-of-domain performance as an optimization problem, as described in Equation 1.
-In practice, this involves repeatedly applying different features within the same foundational model using various hyperparameters $H_{n}^{i}$.
-This process helps us determine the most effective final hyper-parameters $H_{n}$ for each validation center $D_{val}^{n}$.
-We assess the effectiveness of each $H_{n}^{i}$ by evaluating its performance on the OOD validation set $D_{val}^{n}$.
-Next we consider the optimization problem for histopathological tasks.
+The out-of-domain performance of model $M$ would be tested by the features $F_{val}^{n}$ that are preprocessed by the hyper-parameter $H_{n}$ as $y=Infer(H_{n},D_{val}^{n})$, where $y$ is the performance metric.
+We define the histopathological task of identifying the optimal pre-processing for enhanced out-of-domain performance as a hyper-parameter optimization problem, as described in Equation 1, where $\Phi$ is the hyper-parameter space.
+This process helps us determine the most effective final hyper-parameters $H_{n}^{*}$ for each validation center $D_{val}^{n}$.
+$H_{n}^{*}=\underset{H\in\Phi}{\operatorname{arg\,max}}\,\text{Infer}(BAHOP(H,D_{val}^{n}),D_{val}^{n})$ (1)
+4.2 Pre-Processing Problem Modeling
+Next, we consider the optimization problem for histopathological tasks.
 The optimization of $H_{n}$ is actually the optimization of a combination of pre-processing parameters, not a single variable.
@@
 This involves generating a binary mask for the tissue regions by applying a binary threshold to the saturation channel of the downsampled image, following its conversion from RGB to HSV color space.
-Morphological closing and median blurring are employed for smoothing the contours of the detected tissue and in minimizing artifacts.
+Morphological closing and median blurring are employed for smoothing the contours of the detected tissue and minimizing artifacts.
 Subsequently, the approximate contours of both the tissue and the tissue cavities are assessed and filtered based on their area, culminating in the creation of the final segmentation mask.
 We complete set of pre-processing hyperparameters for optimization in histopathological tasks is thus $H_{n}=[x_{1},x_{2},x_{3},x_{4},x_{5},x_{6}]$, where each $x_{i}$ has its own hyper-parameter space.
-Optimization of Fast Hyper-Parameter Search
-Good Pre-Processed WSIs show High Structural Similarity.
-For single-domain inference performance, our experiments demonstrate that pre-processing parameters yielding accurate predictions produce images with similar structures (as illustrated in Figure 5).
-Therefore, we only evaluate the pre-processed Whole Slide Images (WSIs) that closely resemble the optimally pre-processed WSI.
+As illustrated in Fig 4 and the observations discussed in Section 3, the set of pre-processing hyperparameters affect the performance together, varying across the different datasets.
+A large area of tissue is detected as background or holes and then dropped during the pre-processing.
+As illustrated in Fig 4, dropping too many patches decreases performance in Fig.
+4(a) and (c) but increases performance in Fig.
+4(b) where the dataset is changed.
+It is hard and no work to measure how these patches and specific types of cells affect the out-of-domain inference performance in histopathological tasks.
+Thus, we optimize these histopathological tasks by machine learning techniques instead of exploring the impact of the specific type of cells.
+As illustrated in Fig 2, the hyper-parameter optimization for histopathological tasks is non-convex optimization which has many large local optima.
+Therefore, our goal is to search for an optimal hyper-parameter in a short time and with less cost; the optima we search for can just be larger as it can.
+We are not searching for the global optima.
+We just search for an optimal value much better than the default hyper-parameter in a very short time and cost.
+In practice, this involves repeatedly applying different features within the same foundational model using various hyperparameters $H_{n}^{*}$.
+Therefore, we propose a PSNR mechanism as a threshold to advance the speed of hyper-parameter search.
+4.3 Optimization of Fast Hyper-Parameter Search
+4.3.1 Good Pre-Processed WSIs show High Structural Similarity.
+Figure 5: Relationship between PSNR and Accuracy. The red star stands for the reference object. All the hyper-parameters are compared with the reference object to calculate the PSNR.
+For single-domain inference performance, our experiments demonstrate that pre-processing parameters, which lead to similar high accuracy, also generate the pre-processed images with similar structures(as illustrated in Figure 5).
+When the two objects are close in accuracy, their PSNR is always high.
+This phenomenon always exists, no matter how we change the reference object.
+Therefore, we only evaluate the pre-processed Whole Slide Images (WSIs) that closely resemble the current optimally pre-processed WSI in each optimization iteration.
 To quantitatively assess differences between two sets of pre-processing results, we use Peak Signal-to-Noise Ratio (PSNR) to measure the similarity of images generated during patch creation.
-As depicted in Figure 1, our workflow includes creating patches, extracting features, training models, and testing inference data.
+The general histopathological workflow includes creating patches, extracting features, training models, and testing inference data.
 During patch generation, thumbnails are created for visualizing the preprocessing effects.
-Thus, we can determine the similarity between these two sets of pre-processed results by calculating PSNR for these thumbnails.
-PSNR
+Thumbnails are 64x downsampled from original giga-pixel WSIs.
+Thus, we can determine the similarity between these two sets of pre-processed results by calculating PSNR for these thumbnails efficiently.
+4.3.2 Peak Signal-to-Noise Ratio (PSNR)
 Peak Signal-to-Noise Ratio (PSNR) is a common measurement in image processing and digital signal processing.
@@
 For example, the total PSNR of thumbnails produced using the second-best preprocessing parameters should be higher than that of any other set, as illustrated in Figure 5.
-[…]
-In our motivation experiments, which explore the impact of pre-processing parameters on out-of-domain performance, we have already processed sufficient Whole Slide Images (WSIs) with inference results.
-We established the similarity threshold $\tau$ based on previous experimental analyses.
-For our proposed SSAPT method, we use ${H_{train}}$ as the default parameter.
+4.3.3 PSNR-based Basin Hopping for Hyper-Parameter Optimization
+Algorithm 1 PSNR-based Basin Hopping Hyper-Parameter Optimization (BAHOP)
+Input: Default parameter ${H_{train}}$ with result set $\mathcal{R}$, Hyper-parameters space $\Phi$, initial similarity threshold $\tau$, function $Infer$ to evaluate performance, function $S$ to calculate similarity
+Output: Optimal parameters $H^{*}$
+1:  Initialize $H^{*}$ with fixed default parameter ${H_{train}}$
+2:  Generate pre-processed thumbnails $I^{*}$ from $H^{*}$ (for all validation WSIs)
+3:  Generate feature $F^{*}$ from WSIs with $H^{*}$
+4:  $y^{*}\leftarrow Infer(F^{*})$
+5:  $\mathcal{R}\leftarrow(H^{*},y^{*})$
+6:  while optimization times do
+7:     Select $H_{\text{new}}$ by perturbing $H^{*}$ slightly from $\Phi$
+8:     if  $H_{\text{new}}$ not in $\mathcal{R}$ then
+9:        Generate thumbnails $I_{\text{new}}$ with $H_{\text{new}}$
+10:        if $S(I_{\text{new}},I^{*})>\tau$  then
+11:           Extract feature $F_{\text{new}}$      $\triangleright$ Expensive step
+12:           $y_{\text{new}}\leftarrow Infer(F_{\text{new}})$
+13:           $\mathcal{R}\leftarrow\mathcal{R}\cup(H_{\text{new}},y_{\text{new}})$
+14:           if $y_{\text{new}}>y^{*}$  then
+15:              $H^{*},I^{*},y^{*}\leftarrow{H_{\text{new}}},I_{\text{new}},y_{\text{new}}$
+16:           end if
+17:        end if
+18:     end if
+19:  end while
+20:  return  $H^{*}$
+Basin-hopping is a global optimization technique characterized by iterative processes involving random perturbation of coordinates, followed by localized optimization and subsequent evaluation and acceptance or rejection of new coordinates based on minimized function values.
+This method is particularly useful in high-dimensional landscapes.
+However, an inherent limitation of Basin-hopping lies in the necessity to execute random perturbations at a designated point, typically a local minimum.
+These perturbations must be suitably sized - adequately large to escape the current local minimum yet not excessively to devolve into total randomness.
+To address this challenge, we introduce a Peak Signal-to-Noise Ratio (PSNR) threshold within the Basin-hopping optimization process.
+Subsequently, our experimental endeavors focus on establishing a correlation between pre-processing methodologies applied to whole slide images and the PSNR metric.
+The integration of PSNR-based filtering does not fundamentally change the Basin Hopping process itself but rather serves as a heuristic pre-processing step to enhance efficiency and fit histopathological tasks.
+The PSNR threshold is set dynamically and adapts across different datasets.
+The experiments of the effection of each pre-processing parameter are shown in the Appendix.
+From the experiments (as shown in the Appendix), the criteria for selecting the PSNR threshold and dynamically adapting to different datasets is that we perturb the pre-processing parameter – segmentation threshold (related to foreground and background detection).
+Each time, we change the value to plus or minus 1.
+We calculate the PSNR for this comparison as the initial PSNR.
+This process only requires the creation of a patch and quick PSNR calculation in thumbnails.
+Also, we delete the mechanism of temperature and accepted probability from the original Basin hopping, ensuring our BAHOP is designed properly for histopathological tasks.
+For our proposed BAHOP method (as shown in Algorithm 1), we use ${H_{train}}$ as the default parameter.
 Using the current optimal parameter $H^{*}$, we generate a new parameter ${H_{new}}$, create patches, and calculate their PSNR against the optimal thumbnails $I^{*}$.
@@
 Each new parameter set ${H_{new}}$ is added to our historical dataset $R=\{(H_{1},y_{1}),...,(H_{i},y_{i})\}$.
-If ${H_{new}}$ demonstrates improved performance, there is a probability $p$ (as per Eq.(3)) that ${H_{*}}$ will be updated with ${H_{new}}$.
-This process will be further discussed in the following subsection.
-Similarity-based Simulated Annealing for Parameter Tuning
-Algorithm 1 Similarity based Simulated Annealing for Parameter Tuning (SSAPT)
-Input: Default parameter ${H_{train}}$ with result set $\mathcal{R}$, Hyper-parameters space $\Phi$, initial temperature $T_{0}$, minimum temperature $T_{\text{min}}$, cooling rate $\alpha$, similarity threshold $\tau$, function $Infer$ to evaluate performance, function $S$ to calculate similarity, $p$ is the acceptance probability function
-Output: Optimal parameters $H^{*}$
-1:  Initialize $H^{*}$ with fixed default parameter ${H_{train}}$
-2:  Generate pre-processed thumbnail $I^{*}$ from $H^{*}$
-3:  Generate feature $F^{*}$ from WSIs with $H^{*}$
-4:  $y^{*}\leftarrow Infer(F^{*})$
-5:  $\mathcal{R}\leftarrow(H^{*},y^{*})$
-6:  $T\leftarrow T_{0}$
-7:  while $T>T_{\text{min}}$ do
-8:     Select $H_{\text{new}}$ by perturbing $H^{*}$ slightly from $\Phi$
-9:     if  $H_{\text{new}}$ not in $\mathcal{R}$ then
-10:        Generate thumbnail $I_{\text{new}}$ with $H_{\text{new}}$
-11:        $\Delta s\leftarrow S(I_{\text{new}},I^{*})-\tau$
-12:        if $\Delta s>0$  then
-13:           Extract feature $F_{\text{new}}$      $\triangleright$ Expensive step
-14:           $y_{\text{new}}\leftarrow Infer(F_{\text{new}})$
-15:           $\mathcal{R}\leftarrow\mathcal{R}\cup(H_{\text{new}},y_{\text{new}})$
-16:           if $y_{\text{new}}>y^{*}$  then
-17:              $H^{*},I^{*},y^{*}\leftarrow{H_{\text{new}}},I_{\text{new}},y_{\text{new}}\text{ with $p$}$
-18:           end if
-19:        end if
-20:        $T\leftarrow\alpha T$
-21:     end if
-22:  end while
-23:  return  $H^{*}$
-To avoid falling into local optima, we integrate the Simulated Annealing algorithm (Kirkpatrick, Gelatt Jr, and Vecchi 1983) into our method to identify optimal parameters.
-We detail SSAPT in Algorithm 1.
-Our experimental results indicate that creating patches with thumbnails and bypassing most parameter feature extraction processes allows us to refine pre-processing parameters for enhanced inferential performance.
-Our algorithm is designed to optimize pre-processing parameters effectively, ensuring high performance while accelerating the overall process.
-The Simulated Annealing algorithm prevents convergence at local optima by exploring multiple random states near the current solution, akin to simulated annealing itself.
-Occasionally, we enhance solutions by leveraging the current temperature to explore nearby states post-annealing.
-This approach introduces randomness through annealing rules, significantly increasing the likelihood of achieving an optimal solution tailored for histopathological tasks.
-During each iteration, we aim to match the pre-processed results closely with the current optimum.
-We use a similarity threshold $\tau$, related to PSNR as a filter mechanism.
-After patch creation, features are extracted and analyzed only from pre-processed Whole Slide Images (WSIs) that exhibit a PSNR higher than $\tau$.
-WSIs with lower PSNR compared to the optimal pre-processed WSI are excluded from further processing.
-If superior inference performance emerges, updates are made for $H^{*},I^{*},y^{*}$ based on acceptance probability $p$:
-$p(\Delta s)=\exp(\Delta s/T)$ (3)
-Experiments
-Dataset Center ABMIL DSMIL CLAM Bayes-MIL MHIM-DSMIL
-Default Ours Default Ours Default Ours Default Ours Default Ours
-Camelyon17 C0 0.93 0.97 0.77 0.96 0.921 0.95 0.83 0.94 0.87 0.96
-C1 0.87 0.92 0.85 0.89 0.512 0.846 0.35 0.8 0.81 0.87
-C4 0.88 0.93 0.92 0.96 0.874 0.92 0.85 0.92 0.91 0.94
-TCGA-COAD CM 0.75 0.778 0.75 0.778 0.75 0.889 0.861 0.889 0.75 0.801
-D5 0.709 0.774 0.806 0.871 0.742 0.871 0.709 0.806 0.839 0.871
-DM 0.783 0.826 0.869 0.869 0.739 0.739 0.783 0.826 0.826 0.87
-G4 0.963 0.963 0.963 0.963 0.889 0.926 0.926 1 0.925 0.963
-TCGA-BRCA A7 0.902 0.961 0.921 0.941 0.941 0.961 0.882 0.882 0.923 0.961
-AC 0.775 0.8 0.725 0.8 0.75 0.8 0.775 0.775 0.75 0.775
-AR 0.781 0.8 0.765 0.781 0.813 0.828 0.813 0.828 0.75 0.766
-C8 0.907 0.93 0.581 0.813 0.907 0.93 0.86 0.884 0.907 0.93
-Table 2: Experiments in out-of-domain inference performance. We compare the accuracy obtained by default hyper-parameters with the optimal hyper-parameter searched by our SSAPT algorithm.
-Experiments Setup
-Our experiment utilizes the pre-processing pipeline from CLAM.
-The datasets employed in this study include Camelyon 16, Camelyon 17 (Litjens et al.
-2018), TCGA-NSCLC, TCGA-BRCA, and TCGA-COAD (Weinstein et al.
-2013).
-For Camelyon16 and Camelyon17, we focus on normal-tumor classification; for all TCGA datasets, we address subtype tasks.
-In our experiments, we standardize seeds and model hyperparameters such as learning rate and loss weight across similar tasks within the same model framework.
+If ${H_{new}}$ demonstrates improved performance, ${H_{*}}$ will be updated with ${H_{new}}$.
+5 Experiments
+Model TCGA-COAD TCGA-BRCA
+Center Accuracy AUC Center Accuracy AUC
+Default Ours Default Ours Default Ours Default Ours
+ABMIL [12] CM 0.75 0.778 0.828 0.838 A7 0.902 0.961 0.964 0.994
+D5 0.709 0.774 0.64 0.647 AC 0.775 0.805 0.728 0.761
+DM 0.783 0.826 0.732 0.759 AR 0.781 0.816 0.834 0.856
+G4 0.963 0.963 0.98 1 C8 0.907 0.93 0.867 0.908
+DSMIL [15] CM 0.75 0.778 0.848 0.859 A7 0.921 0.941 0.923 0.962
+D5 0.806 0.871 0.5 0.493 AC 0.725 0.805 0.761 0.761
+DM 0.869 0.869 0.777 0.804 AR 0.765 0.781 0.733 0.738
+G4 0.963 0.963 0.96 0.98 C8 0.581 0.813 0.767 0.75
+CLAM [17] CM 0.75 0.889 0.56 0.636 A7 0.941 0.961 0.947 0.974
+D5 0.742 0.871 0.46 0.793 AC 0.75 0.81 0.772 0.761
+DM 0.739 0.739 0.571 0.571 AR 0.813 0.828 0.882 0.878
+G4 0.889 0.926 0.94 0.96 C8 0.907 0.93 0.867 0.875
+BayesMIL [5] CM 0.861 0.889 0.929 0.929 A7 0.882 0.882 0.934 0.955
+D5 0.709 0.806 0.667 0.593 AC 0.775 0.775 0.799 0.821
+DM 0.783 0.826 0.821 0.875 AR 0.813 0.828 0.796 0.791
+G4 0.926 1 0.96 1 C8 0.86 0.884 0.792 0.767
+MHIM- DSMIL  [23] CM 0.75 0.801 0.77 0.798 A7 0.923 0.961 0.957 0.996
+D5 0.839 0.871 0.48 0.493 AC 0.75 0.775 0.611 0.745
+DM 0.826 0.87 0.714 0.786 AR 0.75 0.766 0.709 0.713
+G4 0.925 0.963 0.94 0.98 C8 0.907 0.93 0.742 0.733
+Table 2: Experiments about out-of-domain inference performance in TCGA-COAD and TCGA-BRCA for cancer subtype.
+We compare the accuracy obtained by default hyper-parameters with the optimal hyper-parameter searched by our BAHOP algorithm.
+Model Out-of-Domain(Center 0) Out-of-Domain(Center 1) Out-of-Domain(Center 4)
+Accuracy AUC Accuracy AUC Accuracy AUC
+Def Ours Def Ours Def Ours Def Ours Def Ours Def Ours
+ABMIL [12] 0.93 0.97 0.921 0.965 0.84 0.92 0.833 0.896 0.88 0.93 0.907 0.911
+DSMIL [15] 0.77 0.96 0.88 0.937 0.64 0.89 0.721 0.93 0.92 0.96 0.952 0.947
+CLAM [17] 0.92 0.95 0.919 0.963 0.512 0.846 0.708 0.833 0.874 0.92 0.904 0.941
+Bayes-MIL [5] 0.83 0.94 0.956 0.884 0.35 0.80 0.819 0.881 0.85 0.92 0.931 0.899
+MHIM-DSMIL [23] 0.87 0.96 0.911 0.956 0.77 0.87 0.85 0.921 0.91 0.94 0.949 0.955
+Table 3: Out-of-domain inference performance in Center 0, center1 and Center 4 of Camelyon17 for normal and tumor classification.
+5.1 Experiments Setup
+Our experiment utilizes the pre-processing pipeline from CLAM [17].
+The datasets include Camelyon 16, Camelyon 17 [16], TCGA-NSCLC, TCGA-BRCA, and TCGA-COAD [29].
+For Camelyon16 and Camelyon17, we focus on normal-tumor classification; for all TCGA datasets, we address cancer subtype tasks.
+In our experiments, we fix the seeds and model hyperparameters such as learning rate and loss weight across similar tasks within the same model framework.
 Additional experimental details are provided in the Appendix.
-We employ K-fold, StratifiedKFold or k-fold Monte Carlo cross-validation methods (with k set to 10) to train models and assess out-of-domain performance.
-All results presented in Table 2 and Table 3 represent average accuracies calculated over these 10-fold models.
-Classification Accuracy Validation of SSAPT
-For the Camelyon datasets, we trained our model using 270 whole slide images (WSIs) from Camelyon16 and tested inference performance across various centers in Camelyon17 to simulate out-of-domain scenarios.
-It’s important to note that Camelyon16 data was collected from UMCU and RUMC, while Camelyon17 includes data from CWZ, RST, UMCU, RUMC, and LPON.
-Consequently, only centers 0 (CWZ), 1 (RST), and 4 (LPON) of Camelyon17 represent out-of-domain situations.
+We employ K-fold, Stratified K-Fold or K-fold Monte Carlo cross-validation methods (with k sets to 10) to train models and assess out-of-domain performance.
+All results presented in Table 2, Table 3 and Table 4 represent average accuracy calculated over 10-fold models.
+5.2 Classification Accuracy Validation of BAHOP
+For the Camelyon datasets, we use 270 WSIs from Camelyon 16 to train the model and test inference performance across various centers in Camelyon17 to simulate out-of-domain scenarios.
+It’s important to note that Camelyon16 data was collected from two centers (UMCU and RUMC), while Camelyon17 includes data from five centers (CWZ, RST, UMCU, RUMC, and LPON).
+Consequently, only centers 0 (CWZ), center 1 (RST), and center 4 (LPON) of Camelyon17 represent out-of-domain situations.
+We compare inference performance using features extracted by default parameters (not the worst) and those optimized through our BAHOP algorithm.
+As shown in Table 2 and Table 3, the optimal hyper-parameters identified by our BAHOP algorithm generally surpass the default settings across various models and datasets.
+In specific cases, such as CLAM and BayesMIL at Center 1 of Camelyon 17, performance improved dramatically from 0.512 to 0.847 simply by adjusting the hyper-parameters.
+For TCGA datasets, we divide the entire dataset according to TCGA Tissue Source Site Codes [29].
+We select several centers with sufficient WSIs to form the testing sub-dataset, while the remaining data constitutes the training dataset.
+The out-of-domain inference performance also improves (as shown in Table 2), indicating that our issue is general and our BAHOP provides a universal solution.
+Additionally, we conduct experiments on a larger TCGA-NSCLC dataset that includes more extensive out-of-domain testing (as detailed in Table 4).
+TCGA-NSCLC
+Center ABMIL [12] CLAM [17] Bayes-MIL [5]
+Def. Ours Def. Ours Def. Ours
+21 0.72 0.78 0.56 0.78 0.56 0.67
+22 0.87 0.89 0.78 0.84 0.84 0.84
+43 0.75 0.80 0.60 0.75 0.70 0.80
+49 0.59 0.64 0.69 0.71 0.71 0.76
+55 0.49 0.52 0.66 0.76 0.28 0.33
+77 0.93 0.93 0.91 0.93 0.75 0.80
+97 1.00 1.00 0.91 1.00 0.87 0.96
+Table 4: Comparison of inference accuracy in TCGA-NSCLC with more centers. Only accuracy is reported since each separate center only has one category. Def. stands for default parameter.
+5.3 Comparison of other hyper-parameter optimization
+As illustrated in Table 5, we compare different hyper-parameter optimization (HPO) techniques in one center of Camelyon 17 over 100 iterations.
+Our BAHOP can find the optimal hyper-parameters much quicker compared to alternative HPO methodologies.
+While other HPO techniques demonstrated the capability to achieve commendable accuracy levels, they incurred substantial costs in terms of computational time and resource utilization.
+Table 5: Comparison of other hyper-parameter optimization. SA stands for Simulated Annealing. Lat. stands for Latency (smaller is better) and Mem. stands for Memory.
+Strategy Acc Lat.(min) Mem.(GB)
+Random search 0.845 9600 1250
+Grid search 0.847 9600 1250
+SA [27] 0.846 9600 1250
+Bayes OPT [22] 0.834 9600 1250
+BAHOP (Ours) 0.846 1770 170
+5.4 Efficiency Validation
+Figure 6 shows the running time for optimizing preprocessing parameters 100 times.
+Our PSNR-based Basin Hopping (BAHOP) algorithm for Parameter Tuning uses PSNR to compare each pre-processed WSIs with the best one so far, selecting only those with high similarity for further feature extraction.
+WSIs with low PSNR are skipped, streamlining the process and ensuring that only promising candidates are evaluated further.
+This approach prevents our BAHOP from getting stuck in local optima and saves time by reducing unnecessary feature extractions.
 Figure 6: The running time of pre-processing parameters tuning for 100 times. Our algorithm can save time by skipping a large number of feature extraction.
-We compared the inference performance using features extracted by default parameters (not the worst) and those optimized through our SSAPT algorithm.
-As shown in Table 2, the optimal hyper-parameters identified by our SSAPT algorithm generally surpass the default settings across various models and datasets.
-In specific cases, such as with CLAM and BayesMIL at Center 1, performance improved dramatically from below 0.5 to up to 0.8 simply by adjusting the hyper-parameters.
-Even for robust models like ABMIL, modifying parameters proved effective.
-For TCGA datasets, we divide the entire dataset according to TCGA Tissue Source Site Codes.
-We select several centers with a sufficient number of WSIs to form the testing sub-dataset, while the remaining data constitutes the training dataset.
-The out-of-domain inference performance also improves (as shown in Table 2), indicating that our issue is general and our SSAPT provides a universal solution.
-Additionally, we conduct experiments on a larger TCGA-NSCLC dataset that includes more extensive out-of-domain testing (as detailed in Table 3).
-Dataset Center ABMIL CLAM Bayes-MIL
-Def. Ours Def. Ours Def. Ours
-TCGA- NSCLC 21 0.722 0.778 0.556 0.778 0.556 0.667
-22 0.865 0.892 0.784 0.838 0.838 0.838
-43 0.750 0.800 0.600 0.750 0.700 0.800
-49 0.595 0.643 0.690 0.714 0.714 0.762
-55 0.495 0.516 0.659 0.758 0.275 0.330
-77 0.932 0.932 0.909 0.932 0.750 0.800
-97 1.000 1.000 0.913 1.000 0.869 0.957
-Table 3: Comparison of inference accuracy in TCGA-NSCLC.
-We evaluate the default hyper-parameter with optimal hyper-parameter searched by our SSAPT Algorithm in a larger dataset TCGA-NSCLC with more centers.
-Efficiency Validation
-Figure 4 shows the running time for tuning preprocessing parameters 100 times.
-Our Simulated Annealing algorithm for Parameter Tuning uses PSNR to compare each pre-processed WSI with the best one so far, selecting only those with high similarity for further feature extraction.
-WSIs with low PSNR are skipped, streamlining the process and ensuring that only promising candidates are evaluated further.
-This approach prevents our SSAPT from getting stuck in local optima and saves time by reducing unnecessary feature extractions.
-Ablation Study
-PSNR
-Without the PSNR-based strategy, the algorithm performs a random search.
-It randomly generates pre-processing parameters, creates patches, extracts features, and assesses performance in each iteration without skipping feature extraction (as shown in Table 4).
-Figure 5: The evidence of PSNR-based mechanism works in our fast hyper-parameter search in histopathological tasks.
-We also assess FDD-Simulated Annealing.
-It’s important to note that the Fréchet domain distance (FDD) in this part of the ablation study serves a different role compared to its use in the Discussion section.
-Here, FDD replaces PSNR for comparison purposes.
-Both FDD and PSNR-based mechanisms in this ablation study are employed to constrain the search space for optimal hyper-parameters, rather than searching across the entire hyper-parameter space.
-Table 4: Importance of the PSNR. Results are based on optimizing 100 times in the center1 of Camelyon17. SA stands for Simulated Annealing, Lat. stands for Latency and Mem. stands for Memory.
-Strategy Acc Lat. (min) Mem. (GB)
-None (Baseline) 0.847 9600 1250
-FDD-SA 0.846 5350 650
-PSNR-SA (Ours) 0.846 1770 170
-Simulated Annealing
-Simulated Annealing prevents optimization from getting stuck in local optima.
-As illustrated in Table 5, search algorithms lacking Simulated Annealing repeatedly test hyper-parameters that yield a higher PSNR than the current optimal solution.
-Table 5: Importance of Simulated Annealing (SA). Results based on optimizing 100 times in center1 of Camelyon17.
-Strategy Accuracy Latency (min)
-Without SA 0.829 4119
-With SA 0.846 1770
-Discussion
-Quantifying the Initial Features
-We demonstrate that significant differences exist in the initial features extracted by various pre-processing parameters, which are used as inputs for training models or making inferences.
-These differences arise not only from distribution shifts but also from variations in the number of patches extracted from each WSI, leading to partial information loss.
-For the same WSI, we compare features generated under different hyper-parameter settings: $H_{opt}$, $H_{good}$, and $H_{random}$.
-As shown in Figures 1 and  6, features are similar when their inference performance is comparable.
-However, Figure 6 (Right) highlights distinct differences between $H_{opt}$ and $H_{random}$ within the t-SNE space, underscoring how patch count variations can exacerbate feature representation discrepancies and affect out-of-distribution (OOD) inference performance.
-Figure 6: The visualization of compasion between features.
-Left: The comparison of the feature obtaining high accuracy with the optimal feature.
-Right: The comparison of the feature extracted by random hyper-parameters with the optimal one.
-Note that these features are extracted from the same WSIs.
-Evaluating the Influence of Pre-processing on Model Representations
-We utilize Fréchet domain distance (FDD) (Pocevičiūtė et al.
-2023) to compare features from the penultimate layer and assess how pre-processing parameters influence the model’s representations.
-For each specific out-of-domain testing dataset $D_{val}^{i}$, WSIs undergo pre-processing using $H_{i}$ to generate feature $F_{i}$.
-This feature is then input into the model for inference.
-We extract the penultimate layer feature $P_{i}$ before it reaches the final classifier.
-The same process applies when using optimal pre-processing features identified by our SSAPT algorithm, denoted as $H_{opt}$, resulting in $F_{opt}$.
-For WSIs processed with $H_{i}$, a higher accuracy correlates with a lower FDD compared to those processed with optimal preprocessing ($H_{opt}$), as illustrated in Figure 7.
-Figure 7: Quantifying the Features Impact by Model Visibility.
-SSAPT is designed for histopathological tasks
-Previous hyper-parameter optimization stategies (Falkner, Klein, and Hutter 2018; Choi et al.
-2024) have been developed with a focus on general machine learning models and evaluated on datasets characterized by a large number of images but relatively small image sizes.
-However, the unique characteristics of WSIs should not be ignored.
-WSIs are giga-pixel images that demand significantly higher computational resources.
-Despite advancements in hardware, memory constraints still pose a significant challenge in the digital processing of WSIs in current research (Cifci et al.
-2023; Echle et al.
-2021).
-Additionally, nature images have very different data distribution compared to the whole slide image (Raghu et al.
-2019).
-Therefore, designing hyper-parameter optimization stategy for CPath is necessary.
-The proposed SSAPT algorithm is the first hyper-parameter optimization stategy designed for histopatholocial tasks.
-Conclusion
+5.5 Ablation Study
+Strategy 1: PSNR
+The workflow of inference in histopathological tasks includes creating patches, extracting features, and inferred in a pre-trained model.
+Without the PSNR-based strategy, the algorithm will run the most expensive step – feature extraction, for each iteration.
+Our PSNR machanism can skip feature extraction in many iterations (as shown in Table 6).
+Strategy 2: Basin Hopping
+Basin Hopping prevents optimization from getting stuck in local optima.
+As illustrated in Table 5, search algorithms lacking Basin Hopping repeatedly test hyper-parameters that yield a higher PSNR than the current optimal solution.
+Since other optimizaion techniques have optimized for 100 times, the optimal accuracy is all high but the costs are so huge both in time and disk space.
+Table 6: Ablation Study of the PSNR and Basin Hopping. Results are based on optimizing 100 times in the center1 of Camelyon 17. Lat. stands for Latency (smaller is better) and Mem. stands for Memory.
+PSNR BH Acc Lat. (min) Mem. (GB)
+0.846 9600 1250
+0.845 2553 278
+0.846 1770 170
+6 Conclusion
 Different pre-processing parameters significantly impact feature extraction and model performance in histopathological tasks.
-In this paper, we propose the Similarity-based Simulated Annealing (SSAPT) algorithm for fast parameter tuning, which enhances inference performance on out-of-domain data.
-SSAPT achieves a 5% to 50% accuracy improvement on the Camelyon and multiple TCGA datasets, offering faster hyper-parameter search by reducing feature extraction steps based on WSI characteristics.
+In this paper, we propose the Similarity-based Basin Hopping (BAHOP) algorithm for fast parameter tuning, which enhances inference performance on out-of-domain data.
+BAHOP achieves a 5% to 30% accuracy improvement on the Camelyon and multiple TCGA datasets, offering faster hyper-parameter search by reducing feature extraction steps based on WSI characteristics.